# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 2.4.3. From local to filtered homology
Local symplectic homology spaces are building blocks for the filtered symplectic homology. This principle can be formalized in several ways. For instance, assume that $a$ is the only point of ${\mathcal{S}}(\alpha)$ in an interval $I$. Then
$\sum_{{\mathcal{A}}(x)=a}\dim\operatorname{SH}_{m}(x;{\mathbb{F}})=\dim\operatorname{SH}^{I}_{m}(W;{\mathbb{F}}).$ (2.17)
@@
In a similar vein, we have a spectral sequence
@@
When ${\mathbb{F}}={\mathbb{Q}}$, the flow is non-degenerate and again $I\subset(0,\infty)$, the spectral sequence (2.19), can be reassembled into a complex. Namely, we have a complex generated over ${\mathbb{Q}}$ by good closed orbits with action in $I$, graded by the Conley–Zehnder index and filtered by the action, and such that its homology is $\operatorname{CH}^{I}(W;{\mathbb{Q}})$; see [GG20, Sec. 2.5]. The differential in the complex is completely determined by the same auxiliary data as the one from the construction of the equivariant symplectic or Floer homology; it is functorial in a suitable sense, etc. (Without non-degeneracy condition, the complex is the $E^{1}$-term in (2.19) but then we also need to fix some more additional data to define the differential.) We omit the details of this construction because it is not explicitly used in the paper. Here we only note that this is no longer literally true when ${\mathit{c}har}\,{\mathbb{F}}\neq 0$; for then $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ can, for instance, be infinite.
Finally, let $\zeta_{m}^{-}(a)$ be the number of bars in the barcode of $\operatorname{SH}(W;{\mathbb{F}})$ of degree $m-1$ ending at $a$ and $\zeta^{+}_{m}(a)$ be the number of bars of degree $m$ beginning at $a$. Set $\zeta_{m}(a)=\zeta_{m}^{-}(a)+\zeta_{m}^{+}(a)$. The following observation is nearly obvious.
@@
$\zeta_{m}(a)=\sum_{{\mathcal{A}}(x)=a}\dim\operatorname{SH}_{m}(x).$ (2.21)
@@
It is sufficient to prove the lemma when $a\in{\mathcal{S}}(\alpha)$; for otherwise both sides of (2.21) are zero. Using the background assumption (BA), fix $\epsilon>0$ so small that $a$ is the only point of the intersection $I\cap{\mathcal{S}}(\alpha)$, where $I=(a-\epsilon,a+\epsilon)$. Clearly,
@@ 2.5. The beginning and end maps
For a fixed field ${\mathbb{F}}$, denote by ${\mathcal{B}}$ or ${\mathcal{B}}({\mathbb{F}})$ the collection of bars of the graded persistence module $\operatorname{SH}(W;{\mathbb{F}})$. Assume that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})=0$. Then every bar in ${\mathcal{B}}$ is finite by Theorem 2.3. Furthermore, let ${\mathcal{P}}={\mathcal{P}}({\mathbb{F}})$ be the set of ${\mathbb{F}}$-visible closed Reeb orbits. As in the previous section, let us also formally add to ${\mathcal{P}}$ the set of the “beginning” of the bars of the form $I=(0,a]$ for some $a>0$, denoting the resulting set by ${\mathcal{P}}^{\prime}$. These additional elements $x$ of ${\mathcal{P}}^{\prime}$ are simply in one-to-one correspondence with such bars. By definition, each of such “orbits” $x$ has symplectic homology equal to ${\mathbb{F}}$ and concentrated in $\deg(I)$, and zero action. (We also set $\mu_{-}(x)=\mu_{+}(x):=\deg(x)$ for such $x$.) When $W$ is star-shaped, we are adding just one extra element corresponding to the domain $W$, which we denote by $[W]$. This element has action zero and its supported in degree $n$.
@@
  • $\deg(I)\in\operatorname{supp}\operatorname{SH}(x)$ and $\deg(I)+1\in\operatorname{supp}\operatorname{SH}(y)$.
@@
whenever the equality holds in (2.22). Note also that the second condition of the theorem necessitates that the orbits in ${\mathcal{P}}$ are ${\mathbb{F}}$-visible. We will see later that in the setting of Theorem C and as a consequence of that theorem, the maps $\operatorname{Beg}$ and $\operatorname{End}$ are onto. We do not know if this is true in general. Furthermore, in the setting of Theorem C and again as a consequence of that theorem, both maps are one-to-one. This is certainly not true in general. For instance, when $\dim\operatorname{SH}(x)=3$ and $x$ is the only closed orbit with action ${\mathcal{A}}(x)$, at least one of the pre-images $\operatorname{Beg}^{-1}(x)$ or $\operatorname{End}^{-1}(x)$ must contain more than one point.
@@
$E:=\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}_{m}(x),$
where $\zeta_{m}(a)$ is the bar counting function from Lemma 2.15. To keep track of the orbits, we require in addition the basis elements $e(m,a)_{i}$ to be adapted to the above direct sum decomposition of $E$. Next, let =${\mathcal{B}}=\{I_{1},I_{2},\ldots\}$ be an arbitrary ordering by ${\mathbb{N}}$ of the bars in the symplectic homology barcode ${\mathcal{B}}$. (Note that, by the background assumption (BA), there are countably many bars in ${\mathcal{B}}$.) Let us denote by $\zeta^{i}_{m}(a)$ the restriction of $\zeta_{m}(a)$ to the “truncated” barcodes ${\mathcal{B}}^{i}=\{I_{1},I_{2},\dots,I_{i}\}$. In other words, $\zeta^{i}_{m}(a)$ is the total number of bars of degree $m-1$ ending at $a$ and of degree $m$ beginning at $a$ among the first $i$ bars.
@@
Our requirements on the maps $\operatorname{Beg}$ and $\operatorname{End}$ are deliberately minimal. For instance, with more effort, we could construct these maps so that the sum of the number of bars of degree $m-1$ in $\operatorname{End}^{-1}(x)$ and the number of bars in $\operatorname{Beg}^{-1}(x)$ of degree $m$ is equal to $\dim\operatorname{SH}_{m}(x)$. This would be a refinement of (2.21) in Lemma 2.15. However, in the cases we are interested in, this condition is automatically satisfied as a consequence of (2.21), and we omit this construction.
@@ 3.1.1. Conley–Zehnder index: linear algebra
[…]
Recall that $A\in\operatorname{Sp}(2m)$ is said to be non-degenerate if all eigenvalues of $A$ are different from one, and strongly non-degenerate if all powers $A^{k}$, $k\in{\mathbb{N}}$, are non-degenerate. (Ditto for elements of $\operatorname{\widetilde{Sp}}(2m)$: an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$ is non-degenerate when $\Phi(1)$ is.) We denote the set of non-degenerate symplectic linear maps by $\operatorname{Sp}^{*}(2m)$ and the part of $\operatorname{\widetilde{Sp}}(2m)$ lying over $\operatorname{Sp}^{*}(2m)$ by $\widetilde{\operatorname{Sp}}{}^{*}(2m)$. It is not hard to see that $A=\Phi(1)$ can be connected to a symplectic map with elliptic part equal to $-I$ (if non-trivial) by a path $\Psi$ lying entirely in $\operatorname{Sp}^{*}(2m)$. Concatenating this path with $\Phi$, we obtain a new path $\Phi^{\prime}$. By definition, the Conley–Zehnder index $\mu(\Phi)\in{\mathbb{Z}}$ of $\Phi$ is $\operatorname{\hat{\mu}}(\Phi^{\prime})$. One can show that $\mu(\Phi)$ is well-defined, i.e., independent of $\Psi$. The function $\mu\colon\widetilde{\operatorname{Sp}}{}^{*}(2m)\to{\mathbb{Z}}$ is locally constant, i.e., constant on connected components of $\widetilde{\operatorname{Sp}}{}^{*}(2m)$.
@@
[…]
The upper and lower Conley–Zehnder indices are defined as
$\mu_{+}(\Phi):=\limsup_{\tilde{\Phi}\to\Phi}\mu(\tilde{\Phi})\quad\textrm{and}\quad\mu_{-}(\Phi):=\liminf_{\tilde{\Phi}\to\Phi}\mu(\tilde{\Phi}),$
where in both cases the limit is taken over $\tilde{\Phi}\in\widetilde{\operatorname{Sp}}{}^{*}(2m)$ converging to $\Phi\in\widetilde{\operatorname{Sp}}(2m)$. In fact, $\mu_{+}(\Phi)$ is simply $\max\mu(\tilde{\Phi})$, where $\tilde{\Phi}\in\widetilde{\operatorname{Sp}}{}^{*}(2m)$ is sufficiently close to $\Phi$ in $\operatorname{\widetilde{Sp}}(2m)$; likewise, $\mu_{-}(\Phi)=\min\mu(\tilde{\Phi})$. (In terms of actual paths, rather than their homotopy classes, $\tilde{\Phi}$ can be taken $C^{r}$-close to $\Phi$ for any $r\geq 0$; the resulting definition of $\mu_{\pm}(\Phi)$ is independent of $r$ and equivalent to the one above.) Clearly, $\mu(\Phi)=\mu_{\pm}(\Phi)$ when $\Phi$ is non-degenerate. The indices $\mu_{\pm}$ are the upper semi-continuous and, respectively, lower semi-continuous extensions of $\mu$ from $\widetilde{\operatorname{Sp}}{}^{*}(2m)$ to $\operatorname{\widetilde{Sp}}(2m)$. These maps are of interest to us because they bound the support of the local symplectic homology of an isolated closed orbit; see (2.8) and (2.9)
@@
We say that $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ is a direct sum of $\Phi^{\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ and $\Phi^{\prime\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime\prime})$, where $m=m^{\prime}+m^{\prime\prime}$, and write
$\Phi=\Phi^{\prime}\oplus\Phi^{\prime\prime}\textrm{ when }\Phi(1)=\Phi^{\prime}(1)\oplus\Phi^{\prime\prime}(1)\textrm{ and }\operatorname{\hat{\mu}}(\Phi)=\operatorname{\hat{\mu}}(\Phi^{\prime})+\operatorname{\hat{\mu}}(\Phi^{\prime\prime}).$
@@
We call $\Psi\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ the non-degenerate part of $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ when $\Psi(1)$ is non-degenerate, $\Phi(1)=\Psi(1)\oplus A$, with $m^{\prime}\leq m$ and $A\in\operatorname{Sp}\big{(}2(m-m^{\prime})\big{)}$ totally degenerate (i.e., with all eigenvalues equal to 1), and $\operatorname{\hat{\mu}}(\Psi)=\operatorname{\hat{\mu}}(\Phi)$. When $\Phi$ is itself totally degenerate we say that the non-degenerate part of $\Phi$ is trivial. The non-degenerate part of $\Phi$ is a direct summand.
@@
Consider first a totally degenerate $A\in\operatorname{Sp}(2m)$. Then $A=\exp(JQ)$ where all eigenvalues of $JQ$ are equal zero. The quadratic form $Q$ can be symplectically decomposed into a sum of terms of four types:
@@
These normal forms are taken from [AG, Sec. 2.4] and we find them more convenient to work with than the original Williamson normal forms; see [Wi] and also, e.g., [Ar, App. 6].
[…]
These definitions readily extend to all paths. Namely, every $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ can be written (non-uniquely) a direct sum $\Phi_{0}\oplus\Psi$ where $\Phi_{0}(1)\in\operatorname{Sp}(2m_{0})$ is a totally degenerate and $\Psi(1)\in\operatorname{Sp}(2m_{1})$ is non-degenerate. In particular, $m_{0}$ is the algebraic multiplicity of the eigenvalue 1 of $\Phi(1)$, and $m_{0}+m_{1}=m$. Then we set

### v3

@@ 2.4.3. From local to filtered homology
Local symplectic homology spaces are building blocks for the filtered symplectic homology. This principle can be formalized in several ways. For instance, assume that $a$ is the only point of ${\mathcal{S}}(\alpha)$ in an interval $I=(t,s)$. Then
$\sum_{{\mathcal{A}}(x)=a}\dim\operatorname{SH}_{m}(x;{\mathbb{F}})=\dim\operatorname{SH}^{I}_{m}(W;{\mathbb{F}})$ (2.18)
In fact,
$\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}(x;{\mathbb{F}})=\operatorname{SH}^{I}(W;{\mathbb{F}}).$ (2.19)
by the definition of local symplectic homology. Therefore, we have the long exact sequence
$\ldots\to\operatorname{SH}^{t}_{m}(W;{\mathbb{F}})\to\operatorname{SH}^{s}_{m}(W;{\mathbb{F}})\to\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}_{m}(x;{\mathbb{F}})\to\operatorname{SH}^{t}_{m-1}(W;{\mathbb{F}})\to\ldots.$ (2.20)
@@
In a similar vein, we have analogues of (2.19) and (2.20) for equivariant symplectic homology and a spectral sequence
@@
This can also be easily seen from (2.7) and the equivariant version of (2.20).
@@
When $\mathbb{F}$$\mathbb{Q}$}, the flow is non-degenerate and again $I\subset(0,\infty)$, the spectral sequence (2.22), can be reassembled into a complex. Namely, we have a complex generated over $\mathbb{Q}$} by good closed orbits with action in $I$, graded by the Conley–Zehnder index and filtered by the action, and such that its homology is $\operatorname{CH}^{I}(W;{\mathbb{Q}})$; see [GG20, Sec. 2.5]. The differential in the complex is completely determined by the same auxiliary data as the one from the construction of the equivariant symplectic or Floer homology; it is functorial in a suitable sense, etc. (Without non-degeneracy condition, the complex is the $E^{1}$-term in (2.22) but then we also need to fix more additional data to define the differential.) We omit the details of this construction because it is not explicitly used in the paper. Here we only note that this is no longer literally true when ${\mathit{c}har}\,{\mathbb{F}}\neq 0$; for then $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ can, for instance, be infinite.
Finally, let $\zeta_{m}^{-}(a)$ be the number of bars in the barcode of $\operatorname{SH}(W;{\mathbb{F}})$ of degree m1m-1 ending at $a$ and $\zeta^{+}_{m}(a)$ be the number of bars of degree $m$ beginning at $a$. Set $\zeta_{m}(a)=\zeta_{m}^{-}(a)+\zeta_{m}^{+}(a)$. The following observation is nearly obvious.
@@
$\zeta_{m}(a)=\sum_{{\mathcal{A}}(x)=a}\dim\operatorname{SH}_{m}(x;{\mathbb{F}}).$ (2.24)
@@
Throughout the proof we suppress the ground field $\mathbb{F}$} in the notation. It is sufficient to prove the lemma when $a\in{\mathcal{S}}(\alpha)$ or $a=0$; for otherwise both sides of (2.24) are zero. Using the background assumption (BA), fix $\epsilon>0$ so small that $a$ is the only point of the intersection $I\cap({\mathcal{S}}(\alpha)\cup\{0\})$, where $I=(a-\epsilon,a+\epsilon)$. Clearly,
@@ 2.5. The beginning and end maps
For a fixed field $\mathbb{F}$}, denote by $\mathcal{B}$} or ${\mathcal{B}}({\mathbb{F}})$ the collection of bars of the graded persistence module $\operatorname{SH}(W;{\mathbb{F}})$. Assume that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})=0$. Then every bar in $\mathcal{B}$} is finite by Theorem 2.3. Furthermore, let ${\mathcal{P}}={\mathcal{P}}({\mathbb{F}})$ be the set of $\mathbb{F}$}-visible closed Reeb orbits. As in the previous section, let us also formally add to $\mathcal{P}$} the set of the “beginning” of the bars of the form $I=(0,a]$ for some $a>0$, denoting the resulting set by $\mathcal{P}$$^{\prime}$. These additional elements $x$ of $\mathcal{P}$$^{\prime}$ are simply in one-to-one correspondence with such bars. By definition, each such “orbit” $x$ has symplectic homology equal to $\mathbb{F}$} and concentrated in $\deg(I)$, and zero action. (We also set $\mu_{-}(x)=\mu_{+}(x):=\deg(x)$ for such $x$.) When $W$ is star-shaped, we are adding just one extra element corresponding to the domain $W$, which we denote by $[W]$. This element has action zero and is supported in degree $n$.
@@
  • $\deg(I)\in\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})$ and $\deg(I)+1\in\operatorname{supp}\operatorname{SH}(y;{\mathbb{F}})$.
@@
whenever the equality holds in (2.25). Note also that the second condition of the theorem necessitates that the orbits in $\mathcal{P}$} are $\mathbb{F}$}-visible. We will see later that in the setting of Theorem C and as a consequence of that theorem, the maps $\operatorname{Beg}$ and $\operatorname{End}$ are onto. We do not know if this is true in general. Furthermore, in the setting of Theorem C and again as a consequence of that theorem, both maps are one-to-one. This is certainly not true in general. For instance, when $\dim\operatorname{SH}(x;{\mathbb{F}})=3$ and $x$ is the only closed orbit with action ${\mathcal{A}}(x)$, at least one of the pre-images $\operatorname{Beg}^{-1}(x)$ or $\operatorname{End}^{-1}(x)$ must contain more than one point.
@@
$E:=\bigoplus_{{\mathcal{A}}(x)=a}\operatorname{SH}_{m}(x;{\mathbb{F}}),$
where $\zeta_{m}(a)$ is the bar counting function from Lemma 2.15. To keep track of the orbits, we require in addition the basis elements $e(m,a)_{i}$ to be adapted to the above direct sum decomposition of $E$. Next, let ${\mathcal{B}}=\{I_{1},I_{2},\ldots\}$ be an arbitrary ordering by $\mathbb{N}$} of the bars in the symplectic homology barcode $\mathcal{B}$}. (Note that, by the background assumption (BA), there are countably many bars in $\mathcal{B}$}.) Let us denote by $\zeta^{i}_{m}(a)$ the restriction of $\zeta_{m}(a)$ to the “truncated” barcodes ${\mathcal{B}}^{i}=\{I_{1},I_{2},\dots,I_{i}\}$. In other words, $\zeta^{i}_{m}(a)$ is the total number of bars of degree m1m-1 ending at $a$ and of degree $m$ beginning at $a$ among the first $i$ bars.
@@
Our requirements on the maps $\operatorname{Beg}$ and $\operatorname{End}$ are deliberately minimal. For instance, with more effort, we could construct these maps so that the sum of the number of bars of degree m1m-1 in $\operatorname{End}^{-1}(x)$ and the number of bars in $\operatorname{Beg}^{-1}(x)$ of degree $m$ is equal to $\dim\operatorname{SH}_{m}(x;{\mathbb{F}})$. This would be a refinement of (2.24) in Lemma 2.15. However, in the cases we are interested in, this condition is automatically satisfied as a consequence of (2.24), and we omit this construction.
@@ 3.1.1. Conley–Zehnder index: linear algebra
with $t\in[0,\,1]$, in the self-explanatory notation; see [BG]. The quasimorphism $\operatorname{\hat{\mu}}$ is called the mean index. The continuity requirement holds automatically and is not necessary for the characterization of $\operatorname{\hat{\mu}}$. (This is not obvious). The mean index is also automatically conjugation invariant, as a consequence of the homogeneity, and $\operatorname{\hat{\mu}}$ restricts to an isomorphism $\pi_{1}(\operatorname{Sp}(2m))\to 2{\mathbb{Z}}$. Furthermore, $\operatorname{\hat{\mu}}(\Phi)\in{\mathbb{Z}}$ whenever the elliptic eigenvalues of $\Phi(1)$ are $\pm 1$ or when $\Phi(1)$ is hyperbolic. (Elliptic eigenvalues are by definition unit eigenvalues and the elliptic part of $A\in\operatorname{Sp}(2m)$ is the restriction of $A$ to the sum of the generalized eigenspaces for all elliptic eigenvalues.)
Recall that $A\in\operatorname{Sp}(2m)$ is said to be non-degenerate if all eigenvalues of $A$ are different from one, and strongly non-degenerate if all powers $A^{k}$, $k\in{\mathbb{N}}$, are non-degenerate. (Similarly, an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$ is non-degenerate when $\Phi(1)$ is.) We denote the set of non-degenerate symplectic linear maps by $\operatorname{Sp}^{*}(2m)$ and the part of $\operatorname{\widetilde{Sp}}(2m)$ lying over $\operatorname{Sp}^{*}(2m)$ by $\widetilde{\operatorname{Sp}}{}^{*}(2m)$. It is not hard to see that $A=\Phi(1)$ can be connected to a symplectic map with the elliptic part equal to I-I (if non-trivial) by a path $\Psi$ lying entirely in $\operatorname{Sp}^{*}(2m)$. Concatenating this path with $\Phi$, we obtain a new path $\Phi^{\prime}$. By definition, the Conley–Zehnder index $\mu(\Phi)\in{\mathbb{Z}}$ of $\Phi$ is $\operatorname{\hat{\mu}}(\Phi^{\prime})$. One can show that $\mu(\Phi)$ is well-defined, i.e., independent of $\Psi$. The function $\mu\colon\widetilde{\operatorname{Sp}}{}^{*}(2m)\to{\mathbb{Z}}$ is locally constant, i.e., constant on connected components of $\widetilde{\operatorname{Sp}}{}^{*}(2m)$.
The mean index $\operatorname{\hat{\mu}}(\Phi)$ measures the total rotation angle of certain unit eigenvalues of $\Phi(t)$. For instance, for the path $\Phi(t)=\exp\big(2\pi\sqrt{-1}\lambda t\big)$, $t\in[0,\,1]$, in $\operatorname{Sp}(2)$ we have
$\operatorname{\hat{\mu}}(\Phi)=2\lambda\textrm{ and }\mu(\Phi)=\mathit{sign}(\lambda)\big(2\lfloor|\lambda|\rfloor+1\big)\textrm{ when }\lambda\not\in{\mathbb{Z}},$ (3.1)
where $\mathit{sign}(\lambda)$ stands, predictably, for the sign of $\lambda$.
The upper and lower Conley–Zehnder indices are defined as
@@
Whenever $\Phi^{\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ and $\Phi^{\prime\prime}\in\widetilde{\operatorname{Sp}}(2m^{\prime\prime})$ we have the direct sum $\Phi=\Phi^{\prime}\oplus\Phi^{\prime\prime}\in\widetilde{\operatorname{Sp}}(2m)$, where $m=m^{\prime}+m^{\prime\prime}$, defined in a natural way. This is equivalent to that
@@
We call $\Psi\in\widetilde{\operatorname{Sp}}(2m^{\prime})$ the non-degenerate part of $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ when $\Psi(1)$ is non-degenerate, $\Phi(1)=\Psi(1)\oplus A$, with $m^{\prime}\leq m$ and $A\in\operatorname{Sp}\big(2(m-m^{\prime})\big)$ totally degenerate (i.e., with all eigenvalues equal to 1), and $\operatorname{\hat{\mu}}(\Psi)=\operatorname{\hat{\mu}}(\Phi)$. (The mean index condition is imposed here to make $\Psi$ unique and ensure that $\Psi=\Phi$ when $\Phi$ is non-degenerate.) When $\Phi$ is itself totally degenerate we say that the non-degenerate part of $\Phi$ is trivial. The non-degenerate part of $\Phi$ is a direct summand.
@@
Consider first a totally degenerate symplectic linear map $A\in\operatorname{Sp}(2m)$. Then $A=\exp(JQ)$ where $J$ is the matrix of the standard linear symplectic form on ${\mathbb{R}}^{2n}$ and all eigenvalues of $JQ$ are equal zero. The quadratic form $Q$ can be symplectically decomposed into a sum of terms of four types:
@@
These normal forms, which we will revisit in Section 5.2.1, can be found in [AG, Sec. 2.4] and [CLW, Sect. 2.6] and we find them more convenient to work with than the original Williamson normal forms; see [Wi36] and also, e.g., [Ar, App. 6].
Denoting by $\mathit{sgn}$ the signature, we have $\dim\ker Q_{0}=2$ and $\mathit{sgn}Q_{0}=0$, and $\dim\ker Q_{\pm}=1$ and $\mathit{sgn}Q_{\pm}=\pm 1$. Let $b_{*}(Q)$, where $0,\pm$, be the number of the $Q_{0}$ and $Q_{\pm}$ terms in the decomposition. Let us also set $b_{*}(A):=b_{*}(Q)$ and $\nu_{0}(A):=\nu_{0}(Q)$, and $b_{*}(\Phi):=b_{*}(A)$ when $A=\Phi(1)$ is totally degenerate. These are symplectic invariants of $Q$ and $A$ and $\Phi$, and the geometric multiplicity of the eigenvalue $1$ is $2(b_{0}+\nu_{0})+b_{+}+b_{-}$.
These definitions readily extend to all paths. Namely, every $\Phi\in\widetilde{\operatorname{Sp}}(2m)$ can be written (non-uniquely) as a direct sum $\Phi_{0}\oplus\Psi$ where $\Phi_{0}(1)\in\operatorname{Sp}(2m_{0})$ is totally degenerate and $\Psi(1)\in\operatorname{Sp}(2m_{1})$ is non-degenerate. In particular, $m_{0}$ is the algebraic multiplicity of the eigenvalue 1 of $\Phi(1)$, and $m_{0}+m_{1}=m$. Then we set